OpenStudy (anonymous):

CAN SOME PLESE HELP ME~!!!!!! of the following is not an equivalent form of the compound inequality x - 1 < -11 or x + 12 > 11 x < -10 or x > -1 A number line with an open circle on -10, shading to the left, and an open circle on -1, shading to the right. x > -1 or x < -10 A number line with an open circle on -10, an open circle on -1, and shading in between.

OpenStudy (anonymous):

@kelliegirl33 @Ezizio @Frostbite @Tahamohammed @GingerNerd @genius12

OpenStudy (anonymous):

I just did this...lol.....last one is incorrect

OpenStudy (anonymous):

x + 16 less than or greater to 11 and x + 16 greater than or equal to 3 11 greater than or equal to x + 16 greater than or equal to 3 -5 greater than or equal to x greater than or equal to -13 A number line with a closed circle on -13, a closed circle on -5, and shading in between. A number line with a closed circle on -13, shading to the left, and a closed circle on -5, shading to the right

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 16 less than or greater to 11 and x + 16 greater than or equal to 3

OpenStudy (anonymous):

x + 16 <= 11 x + 16 >= 3 x <= 11 - 16 x >= 3 - 16 x <= - 5 x >= - 13 Last one is incorrect

OpenStudy (anonymous):

ok this is the last one Choose the correct description of the graph of the compound inequality x - 2 > -3 and 3x less than or greater to 12 A number line with an open circle on -1, shading to the left, and a closed circle on 4, shading to the right. A number line with a closed circle on -1, shading to the left, and an open circle on 4, shading to the right. A number line with an open circle on -1, a closed circle on 4, and shading in between. A number line with a closed circle on -1, an open circle on 4, and shading in between.

OpenStudy (anonymous):

x - 2 > -3 3x <= 12 x > -3 + 2 x <= 12/3 x > -1 x <= 4 3rd one is correct

OpenStudy (anonymous):

THANKKKKKKKKS! <3333

OpenStudy (anonymous):

no problem :)